Isolation Inc. purchased a machine for $450,000. The machine has a $2,000 residual value and is expected to produce 1,500,000 units.
What is the depreciation per unit?
If in year one the machine produces 225,000 units, what would depreciation be?
If in year two the machine produces 1,000,000 units, what would depreciation be?
Answer #1

a.depreciation per unit=(Cost-Residual value)/Useful Life

=(450,000-2000)/1,500,000=$0.29867 per unit(Approx)

b.Depreciation =0.29867*225,000

=$67200

c.Depreciation =0.29867*1,000,000

=$298666.67(Approx).